OpenStudy (anonymous):

Find the equation in point-slope form passing through the points (3, -12) and (-4,2).

jimthompson5910 (jim_thompson5910):

First find the slope m = (2-(-12))/(-4-3) = 14/(-7) = -2 So m = -2 Now plug x=3, y = -12 and m=-2 into the equation y - y1 = m(x-x1) to get y-(-12) = -2(x-3) and simplify y + 12 = -2(x-3)

OpenStudy (radar):

Since you were given two points you can use the: y2 - y1 m = --------- x2-x1 m=(2-(-12))/(-4-(3))=14/-7=-2 slope is -2 or m=-1 It appears that jim_thompson beat me to the finish. lol

OpenStudy (anonymous):

12 + y = -2(x + -3) Reorder the terms: 12 + y = -2(-3 + x) 12 + y = (-3 * -2 + x * -2) 12 + y = (6 + -2x)

OpenStudy (radar):

The equation is normally placed in the form of: y=-2x-6 (slope intercept form 2x+y+6=0 standard form

OpenStudy (anonymous):

ok i think i get it

OpenStudy (radar):

Standard form may also be: 2x+y=-6